OpenStudy (barrelracer011):

Choose the correct simplification of the expression the quantity 4 times x all over y all to the second power. 4 times x to the second power all over y to the second power 8 times x to the second power all over y to the second power 16 times x to the second power all over y to the second power 16 times x to the second power all over y

OpenStudy (barrelracer011):

@amistre64

OpenStudy (johnweldon1993):

\[\large (\frac{4x}{y})^2\] Like that?

OpenStudy (barrelracer011):

Yes I think the answer it either A or B but im not sure

OpenStudy (johnweldon1993):

Not quite...since \[\large (\frac{4x}{y})^2 = \frac{(4x)^2}{y^2}\]

OpenStudy (barrelracer011):

So I would have to do 4x times 4x and get 16x over y^2

OpenStudy (barrelracer011):

So It would be C

OpenStudy (johnweldon1993):

And we know that 4x is actually just written as \(\large 4 \times x\) So really we have \[\large \frac{(4 \times x)^2}{y^2} \rightarrow \frac{(4^2 \times x^2)}{y^2}\]

OpenStudy (johnweldon1993):

C would be correct!

OpenStudy (barrelracer011):

Thank you soo much! Can I mention you in more of these questions if I have any?

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (barrelracer011):

OK thanks!!!